OpenStudy (anonymous):

Find the area bounded by the parabola y=(1/2)x^2 and the hyperbola y^2-x^2=8. state your answer exactly.

OpenStudy (anonymous):

If you solve for \(y\) in the hyperbola, you get: \[ y=\pm \sqrt{x^2+8} \]

OpenStudy (anonymous):

Both equations are symmetric across the \(y\) axis, you can can integrate one section and double the result.

OpenStudy (anonymous):

so then \[(1/2)x^2 \pm \sqrt{x^2 +8} = 0\]

OpenStudy (anonymous):

You'll want to integrate with respect to \(y\) too, since this is a \(y\) simple area.

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

I don't need intercepts?

OpenStudy (anonymous):

You need to find points of intersection to find limits of integration.

OpenStudy (anonymous):

Hold on, graph it first.

OpenStudy (anonymous):

Okay so it looks like this: